# Machine Learning Clifford invariants of ADE Coxeter elements

### v1

@@ 1. Introduction
Recent work has sparked great interest in Clifford geometric invariants of linear transformations [52, 53, 63, 1, 33], originally proposed in [45]. Linear transformation invariants are traditionally exemplified by the determinant and trace, which appear in the highest and lowest coefficients of the characteristic polynomial. Typically such linear transformations are described by matrices; however, in Clifford algebras one has the alternative to implement orthogonal transformations via versors. In Clifford algebras, algebraic objects have a clearer geometric interpretation than in the standard matrix approach. There is a systematic way of calculating multivector invariants of linear transformations via what are called ‘simplicial derivatives’, which we will introduce further in the next section. These Clifford geometric invariants are then systematically related to geometric invariant spaces of the linear transformation and the coefficients in the characteristic polynomial and Cayley-Hamilton theorem [^1] The decomposition of a linear transformation into orthogonal eigenspaces is also related to some interesting recent work by [61].. This serves as motivation to study this type of new geometric invariant of linear transformations.
@@
In previous work, the authors have established a paradigm for experimental mathematics: first, using computational algebra techniques and high performance computing (HPC) one can generate a dataset of algebraic data; this dataset can then be mined by applying the standard data science toolkit, in order to find patterns that were not obvious from an analytical perspective [32, 25, 39, 42, 41, 38]. Mathematicians often calculate examples of interest by hand to formulate or test hypotheses. Essentially, this computational algebra approach automates and scales up such an approach, and turns such examples into ‘big data’ [^3] Clifford algebra multivector computations can easily be performed in such a python HPC setup using the galgebra package [20].. One can either calculate a very large number of examples and analyse these statistically via ‘big data’ data science; or in other cases of interest it may be possible to calculate all the cases exhaustively and analyse the patterns that emerge, which can help with hypothesis formulation and theorem proving.
@@ 2. Background
[…]
(knoten1) edge (knoten2); (knoten2) edge (knoten3); (knoten3) edge (knoten4); (knoten4) edge (knoten5); (knoten5) edge (knoten6); (knoten6) edge (knoten7); (knoten7) edge (knoten8);
[…]
(knoten1) edge (knoten2); (knoten2) edge (knoten3); (knoten3) edge (knoten4); (knoten4) edge (knoten5); (knoten5) edge (knoten6); (knoten6) edge (knoten7); (knoten6) edge (knoten8);
[…]
(knoten1) edge (knoten2); (knoten2) edge (knoten3); (knoten3) edge (knoten4); (knoten4) edge (knoten5); (knoten5) edge (knoten6); (knoten6) edge (knoten7); (knoten5) edge (knoten8);
@@
Since we are considering an even orthogonal transformation in an 8-dimensional space we get some interesting structure in these invariants (see Table 1): firstly, we note that only even multivectors occur (in principle, this allows us to reduce the length of the 256-dimensional multivectors by half). Secondly, the lowest order invariant only has a scalar part (trivially), the next picks up a bivector term, the next one a quadrivector term, the next a sextuvector, till finally $\mathop{\mathrm{Inv}}_{4}$ (generically) has a pseudoscalar term. Then it decreases again. In fact, thirdly, in our case we have a certain ‘mirror symmetry’, where the top half in the Table is equal to the bottom half, though this is not generally the case. In fact, all these pieces, which we could denote by $\mathop{\mathrm{Inv}}_{r}^{k}$ are separately invariant under the Coxeter versor: $\tilde{W}\mathop{\mathrm{Inv}}_{r}^{k}W=\mathop{\mathrm{Inv}}_{r}^{k}$. So these $\mathop{\mathrm{Inv}}_{r}^{k}$ are eigenmultivectors of the Coxeter element of grade $k$, but they do not have to be $k$-blades (i.e. be able to be written as the outer product of $k$ vectors [^9] Something also noticed in the example in [53].).
@@
Invariants by Order scalar bivector quadrivector sextuvector pseudoscalar
@@ 3. Datasets
We choose dimension 8 because of the following compromise: $8!=40320$ gives us something resembling ‘big data’ which is accessible to data science techniques, whilst being computationally tractable [^10] With the caveat that there is degeneracy in the permutations leading to the same or similar Coxeter elements and thus invariants, reducing the true number of different output vectors.. It is also the last dimension in which there are three simply-laced root systems, with the exceptional $E_{8}$ adding some variety to the $A_{n}$ and $D_{n}$ families that exist in arbitrary dimensions. So we select $A_{8}$, $D_{8}$ and $E_{8}$, as this gives us scope for three-way (ternary) classification tasks and ADE patterns are of course of wider interest.
@@ 3.1. Data Generation
The computational algebra approach followed here used python with the galgebra package for multivector computations [20]. Exploratory analysis for single permutations was performed in Jupyter notebooks but once parallelised the computations were run on clusters at Queen Mary, University of London, City, University of London, and University of Leeds. Data and Code can be found on GitHub. We performed computations both in a Euclidean basis, which is a bit more straightforward, and the basis of simple roots via the multivector basis that it induces, which is more meaningful geometrically and less dependent on the choice of simple roots in the Euclidean basis. According to our earlier discussion around Table 1 we can also extract different grades of these invariants (e.g. scalar, bivector etc), which we refer to as ‘subinvariants’ $\mathop{\mathrm{Inv}}_{r}^{k}$, from the full set (SOCM).
@@ 3.2. Frequency Analysis
These exhaustive computational algebra calculations already show interesting results. On the highest level, we compare components of SOCMs and find that, although there are 40320 SOCMs we can construct, only 128 are distinct, and this is the same number for each of the algebras. We can think of these as ‘classes’ of SOCMs with the same components. Each class has a certain number of representatives [^12] They have different order of roots in a permutation encoding Coxeter versor. in it which we call frequency. Although the number of classes for each of the algebras is the same, frequencies of individual classes differ, see Figure 2.
@@
Figure 2. Sorted multiplicities of the 128 unique SOCMs, for each root system considered: $A_{8}$, $D_{8}$, $E_{8}$ respectively. $A_{8}$ is mostly quadruplets, $E_{8}$ mostly doublets and $D_{8}$ half and half.
@@
  • The frequency values comes in groups of 2, 4 or 8 classes, which all have the same frequency. We call these groups of classes with the same frequency as Doublets, Quadruplets and Octuplets, respectively;
@@
Invariant Order scalar bivector quadrivector sextuvector pseudoscalar
@@
Invariant Order scalar bivector quadrivector sextuvector pseudoscalar
@@
Invariant Order scalar bivector quadrivector sextuvector pseudoscalar
@@
Invariant Order scalar bivector quadrivector sextuvector pseudoscalar
@@
Invariant Order scalar bivector quadrivector sextuvector pseudoscalar
@@
Another interesting thing to look at is the frequencies of invariants and subinvariants with the identification of objects that differ up to an overall minus sign. We find that this modification does not alter the frequencies of the full invariants, however, it does change the frequencies of subinvariants, see Tables 6, 6, 6. We see that frequencies for bivector and sextuvector (and pseudoscalar for $A_{8}$) subinvariants for $A_{8}$ and $D_{8}$ are halved, meaning that half of these subinvariants differ from the other half by a minus sign. At the same time, scalar and quadrivector subinvariants are unchanged. The picture is different for $E_{8}$, where bivector, quadrivector and sextuvector frequencies change non-trivially under sign identification.
@@ 4. Machine Learning
ML as a field subdivides into three core categories, the first being supervised learning where many-parameter functions are fitted to large datasets of (input, output) pairs. The second is unsupervised learning; here there is no respective output data, and traditional data analysis methods for feature extraction and clustering are applied directly to the datasets. The third is reinforcement learning, a selection of techniques for training an agent to optimally search a space of potential solutions to a problem – less relevant to this study; focus is thus put on supervised and unsupervised techniques.
@@ 4.1.1. Binary Classification of Invariants: real vs fake data
The dataset we are using consists of 3 datasets, one for each of the algebras $A_{8}$/$E_{8}$/$D_{8}$, with 40320 entries and 2304 components in each entry. To further enlarge the amount of data, we generated a fake dataset for each of the $A_{8}$/$E_{8}$/$D_{8}$ algebra. In the first iteration, ‘fake’ datasets were generated by constructing empirical distributions for each of the 2304 components from all available invariants for the corresponding algebra and then sampling from this distribution to create 40,000 unique elements in new datasets. One can notice that in the ‘real’ data, the number of zeros in each entry is constant and specific for different algebras 1919191942 for $E_{8}$, 2083 for $D_{8}$ and 1805 for $A_{8}$.. We implemented this feature in fake datasets we used by eliminating fake data entries which do not satisfy this condition.
@@
We tried NNs with architecture varying from 1 hidden layer with 32 units to 2 hidden layers with 256 units in each. In all of them, the ReLU activation function was used. During training, we used the Adam optimizer (with a learning rate of 0.001) to optimize the log-loss function. The train and test datasets represented an 80% / 20% split of the total data.
Initially, we trained three different NNs to distinguish one of the $A_{8}/D_{8}/E_{8}$ invariants from others and fake invariants using all datasets. For all three of them, the prediction on test datasets was 100% accurate. This result may be expected because, as described in section 3.2, there are only 128 unique components for each of the ADE datasets, leading to a large repetition of many invariants (although with unequal frequencies). Hence, there is a high chance that in a randomly chosen training subset, we will find all 128 unique invariants. This would make the test stage biased, as it likely contains repetitions of the training data. Effectively, these NNs are learning to reproduce this dataset of invariants perfectly, but would not be able to generalise beyond it.
A more meaningful problem is to remove degeneracy in datasets for ‘real’ $A_{8}/D_{8}/E_{8}$ invariants, leaving just 128 elements in each. Then we can mix original and fake invariants, which ensures that we have some real invariants in the training set and others in the test set only. The proportion of data in training and test datasets was again set to 80% and 20%. However, this makes the whole dataset skewed as there are around 40000 fake invariants and only $3\times 128$ real invariants. We kept data unbalanced in the training set, but in the test set, to make it easier to interpret results, we cut the number of fake invariants to be the same as the number of real invariants, meaning we had $3\times\{128/k$ real and $128/k$ fake} invariants.
In this setup, we tried NN with architectures varying from 1 hidden layer with 32 units to 2 hidden layers with 256 units in each. The best performance was demonstrated by 1 hidden layer 64, 128 and 256 units NNs with accuracies in the range of 0.90-0.92. From this one might speculate that there should be some relatively simple invariant quantity (similar to the genus of a surface) that was learned by the NNs to distinguish real invariants coming from different algebras and fake invariants.
@@ 4.1.2. Regressing Invariants from Permutations
In this paper, we used dense NNs coded in python where the input is the (one-hot encoded [6]) permutation and the output is the coefficients of the invariants. For clearer results, we first partitioned each dataset for $A_{8},D_{8},E_{8}$ into 9 subdatasets for each of the 9 invariant orders, and then look at the coefficients of both the full invariant and each subinvariant for each order subdataset, i.e., the scalar, bivector, quadrivector, sextuvector, and pseudoscalar for each invariant order $\mathop{\mathrm{Inv}}_{0}$-$\mathop{\mathrm{Inv}}_{8}$ (SOCM). The NN model includes four dense layers of 256 units with ReLU activation function. In training, we used the Adam optimizer (with learning rate 0.001) to minimise a mean-squared error loss. 5-fold cross validation was also used, with test data subset being $20\%$ of the full dataset. To calculate accuracy, predictions were rounded to the nearest integer, and a prediction was considered correct if all of its coefficients were predicted correctly after rounding.
Our ML results are summarised in Tables 9 to 9 for the $A_{8},D_{8},E_{8}$ data in the simple root basis. The results show near-perfect prediction of all invariants and subinvariants across all algebras. The trivial scalar invariants are unsurprisingly all learnt perfectly, but in many other cases there is perfect learning also. The lowest performance occurs for the sextuvectors, where there is less data to learn from. These results indicate that the NNs are capable of well approximating the complicated algorithm carried out to compute these invariants, as well as accommodating for the basis permutations.
@@
Acc($\mathop{\mathrm{Inv}}_{i}$)
Acc($\mathop{\mathrm{Inv}}_{i}^{0}$)
Acc($\mathop{\mathrm{Inv}}_{i}^{2}$)
Acc($\mathop{\mathrm{Inv}}_{i}^{4}$)
Acc($\mathop{\mathrm{Inv}}_{i}^{6}$)
Acc($\mathop{\mathrm{Inv}}_{i}^{8}$)
$\mathop{\mathrm{Inv}}_{0}$
$1.0000$
$1.0000$
$\mathop{\mathrm{Inv}}_{1}$
$1.0000$
$1.0000$
$0.9996$
$\mathop{\mathrm{Inv}}_{2}$
$0.9996$
$1.0000$
$0.9999$
$0.9999$
$\mathop{\mathrm{Inv}}_{3}$
$0.9980$
$1.0000$
$0.9973$
$0.9480$
$0.9999$
$\mathop{\mathrm{Inv}}_{4}$
$0.9958$
$1.0000$
$0.9999$
$0.9999$
$0.9117$
$0.9596$
$\mathop{\mathrm{Inv}}_{5}$
$0.9986$
$1.0000$
$0.9995$
$0.9999$
$1.0000$
$\mathop{\mathrm{Inv}}_{6}$
$1.0000$
$1.0000$
$0.9948$
$0.9999$
$\mathop{\mathrm{Inv}}_{7}$
$0.9999$
$1.0000$
$1.0000$
$\mathop{\mathrm{Inv}}_{8}$
$1.0000$
$1.0000$
Acc($\mathop{\mathrm{Inv}}_{i}$)
Acc($\mathop{\mathrm{Inv}}_{i}^{0}$)
Acc($\mathop{\mathrm{Inv}}_{i}^{2}$)
Acc($\mathop{\mathrm{Inv}}_{i}^{4}$)
Acc($\mathop{\mathrm{Inv}}_{i}^{6}$)
Acc($\mathop{\mathrm{Inv}}_{i}^{8}$)
$\mathop{\mathrm{Inv}}_{0}$
$1.0000$
$1.0000$
$\mathop{\mathrm{Inv}}_{1}$
$1.0000$
$1.0000$
$0.9955$
$\mathop{\mathrm{Inv}}_{2}$
$1.0000$
$1.0000$
$0.9912$
$0.9999$
$\mathop{\mathrm{Inv}}_{3}$
$0.9993$
$1.0000$
$0.9995$
$0.9999$
$1.0000$
$\mathop{\mathrm{Inv}}_{4}$
$0.9995$
$1.0000$
$0.9988$
$0.9891$
$0.9998$
$1.0000$
$\mathop{\mathrm{Inv}}_{5}$
$0.9995$
$1.0000$
$0.9986$
$1.0000$
$1.0000$
$\mathop{\mathrm{Inv}}_{6}$
$1.0000$
$1.0000$
$1.0000$
$1.0000$
$\mathop{\mathrm{Inv}}_{7}$
$1.0000$
$1.0000$
$0.9999$
$\mathop{\mathrm{Inv}}_{8}$
$1.0000$
$1.0000$
Acc($\mathop{\mathrm{Inv}}_{i}$)
Acc($\mathop{\mathrm{Inv}}_{i}^{0}$)
Acc($\mathop{\mathrm{Inv}}_{i}^{2}$)
Acc($\mathop{\mathrm{Inv}}_{i}^{4}$)
Acc($\mathop{\mathrm{Inv}}_{i}^{6}$)
Acc($\mathop{\mathrm{Inv}}_{i}^{8}$)
$\mathop{\mathrm{Inv}}_{0}$
$1.0000$
$1.0000$
$\mathop{\mathrm{Inv}}_{1}$
$1.0000$
$1.0000$
$1.0000$
$\mathop{\mathrm{Inv}}_{2}$
$0.9999$
$1.0000$
$1.0000$
$1.0000$
$\mathop{\mathrm{Inv}}_{3}$
$0.9994$
$1.0000$
$0.9906$
$0.9999$
$0.9891$
$\mathop{\mathrm{Inv}}_{4}$
$0.9969$
$1.0000$
$1.0000$
$0.9963$
$0.9793$
$0.9223$
$\mathop{\mathrm{Inv}}_{5}$
$0.9994$
$1.0000$
$0.9996$
$0.9999$
$0.9990$
$\mathop{\mathrm{Inv}}_{6}$
$1.0000$
$1.0000$
$1.0000$
$1.0000$
$\mathop{\mathrm{Inv}}_{7}$
$1.0000$
$1.0000$
$0.9998$
$\mathop{\mathrm{Inv}}_{8}$
$1.0000$
$1.0000$

### v2

@@ 1. Introduction
Great interest in Clifford geometric invariants of linear transformations, originally proposed in [45], was sparked in recent work from a practical [53, 52] and theoretical [52, 63, 1, 33] point of view. Orthogonal transformations, such as rotations, and their invariants are important in engineering, e.g. moving cameras, robots etc. The types of transformations we are looking at in this work are also rotations, particularly interesting because of their symmetry structures. Linear transformation invariants are traditionally exemplified by the determinant and trace, which appear in the highest and lowest coefficients of the characteristic polynomial. Typically such linear transformations are described by matrices; however, in Clifford algebras one has the alternative to implement orthogonal transformations via versors. In Clifford algebras, algebraic objects have a clearer geometric interpretation than in the standard matrix approach. There is a systematic way of calculating multivector invariants of linear transformations via what are called ‘simplicial derivatives’, which we will introduce further in the next section. These Clifford geometric invariants are then systematically related to geometric invariant spaces of the linear transformation and the coefficients in the characteristic polynomial and Cayley-Hamilton theorem [^1] The decomposition of a linear transformation into orthogonal eigenspaces is also related to some interesting recent work by [61].. This serves as motivation to study this type of new geometric invariant of linear transformations.
@@
In previous work, the authors have established a paradigm for experimental mathematics: first, using computational algebra techniques and high-performance computing (HPC) one can generate a dataset of algebraic data; this dataset can then be mined by applying the standard data science toolkit, in order to find patterns that were not obvious from an analytical perspective [32, 25, 39, 42, 41, 38]. Mathematicians often calculate examples of interest by hand to formulate or test hypotheses. Essentially, this computational algebra approach automates and scales up such an approach, and turns the problem into a ‘data analysis’ task  [^3] Clifford algebra multivector computations can easily be performed in such a python HPC setup using the galgebra package [20].. One can either calculate a very large number of examples and analyse these statistically via ‘data analysis’ ; or in other cases of interest, it may be possible to calculate all the cases exhaustively and analyse the patterns that emerge, which can help with hypothesis formulation and theorem proving.
@@ 2. Background
$\alpha_{1}$$\alpha_{2}$$\alpha_{3}$$\alpha_{4}$$\alpha_{5}$$\alpha_{6}$$\alpha_{7}$$\alpha_{8}$
$\alpha_{1}$$\alpha_{2}$$\alpha_{3}$$\alpha_{4}$$\alpha_{5}$$\alpha_{6}$$\alpha_{7}$$\alpha_{8}$
$\alpha_{1}$$\alpha_{2}$$\alpha_{3}$$\alpha_{4}$$\alpha_{5}$$\alpha_{6}$$\alpha_{7}$$\alpha_{8}$
@@
Since we are considering an even orthogonal transformation in an 8-dimensional space we get some interesting structure in these invariants (see Table 1): firstly, we note that only even multivectors occur (in principle, this allows us to reduce the length of the 256-dimensional multivectors by half). Secondly, the lowest order invariant only has a scalar part (trivially), the next picks up a bivector term, the next one a quadrivector term, the next a sextivector, till finally $\mathop{\mathrm{Inv}}_{4}$ (generically) has a pseudoscalar term. Then it decreases again. In fact, thirdly, in our case we have a certain ‘mirror symmetry’, where the top half in the Table is equal to the bottom half, though this is not generally the case. In fact, all these pieces, which we could denote by $\mathop{\mathrm{Inv}}_{r}^{k}$ are separately invariant under the Coxeter versor: $\tilde{W}\mathop{\mathrm{Inv}}_{r}^{k}W=\mathop{\mathrm{Inv}}_{r}^{k}$. So these $\mathop{\mathrm{Inv}}_{r}^{k}$ are eigenmultivectors of the Coxeter element of grade $k$, but they do not have to be $k$-blades (i.e. be able to be written as the outer product of $k$ vectors [^9] Something also noticed in the example in [53].).
@@
Invariants by Order scalar bivector quadrivector sextivector pseudoscalar
@@ 3. Datasets
We choose dimension 8 because of the following compromise: $8!=40320$ gives us something resembling ‘big data’ which is accessible to data science techniques, whilst being computationally tractable [^10] With the caveat that there is degeneracy in the permutations leading to the same or similar Coxeter elements and thus invariants, reducing the true number of different output vectors. Although it was not obvious from the beginning, especially for $E_{9}$ and $D_{8}$. . It is also the last dimension in which there are three simply-laced root systems, with the exceptional $E_{8}$ adding some variety to the $A_{n}$ and $D_{n}$ families that exist in arbitrary dimensions. So we select $A_{8}$, $D_{8}$ and $E_{8}$, as this gives us scope for three-way (ternary) classification tasks and ADE patterns are of course of wider interest.
@@ 3.1. Data Generation
The computational algebra approach followed here used python with the galgebra package for multivector computations [20]. Exploratory analysis for single permutations was performed in Jupyter notebooks but once parallelised the computations were run on clusters at Queen Mary, University of London, City, University of London, and University of Leeds. Data and Code can be found on GitHub. We performed computations both in an Euclidean basis121212 Roots in the root system are often defined as columns of components in the Euclidean orthonormal basis in some higher dimensional space. However, since the set of simple roots can generate the root system via addition, we can also take simple roots as a basis, although not orthonormal. While this may seem more complicated, it is more meaningful geometrically because everything we compute in geometric algebra using simple roots can be eventually expressed in the basis of simple roots. , which is a bit more straightforward, and the basis of simple roots via the multivector basis that it induces, which is more meaningful geometrically and less dependent on the choice of simple roots in the Euclidean basis. According to our earlier discussion around Table 1 we can also extract different grades of these invariants (e.g. scalar, bivector etc), which we refer to as ‘subinvariants’ $\mathop{\mathrm{Inv}}_{r}^{k}$, from the full set (SOCM).
@@ 3.2. Frequency Analysis
These exhaustive computational algebra calculations already show interesting results. On the highest level, we compare components of SOCMs and find that, although there are 40320 SOCMs we can construct, only 128 are distinct, and this is the same number for each of the algebras. We can think of these as ‘classes’ of SOCMs with the same components. Each class has a certain number of representatives [^13] They have different order of roots in a permutation encoding Coxeter versor. in it which we call frequency. The frequency of classes, which come in groups of 2, 4 or 8 and all have the same value, we call Doublets, Quadruplets and Octuplets, respectively. Although the number of classes for each of the algebras is the same, frequencies of individual classes differ, see Figure 2.
@@
Figure 2. Sorted multiplicities of the 128 unique SOCMs, for each root system considered: $A_{8}$, $D_{8}$, $E_{8}$ respectively. $A_{8}$ is mostly quadruplets, $E_{8}$ mostly doublets and $D_{8}$ half and half. See GitHub for the full list of values.
@@
  • All the frequency values come in Doublets, Quadruplets and Octuplets;
@@
Invariant Order scalar bivector quadrivector sextivector pseudoscalar
@@
Invariant Order scalar bivector quadrivector sextivector pseudoscalar
@@
Invariant Order scalar bivector quadrivector sextivector pseudoscalar
@@
Invariant Order scalar bivector quadrivector sextivector pseudoscalar
@@
Invariant Order scalar bivector quadrivector sextivector pseudoscalar
@@
Another interesting thing to look at is the frequencies of invariants and subinvariants with the identification of objects that differ up to an overall minus sign. We find that this modification does not alter the frequencies of the full invariants, however, it does change the frequencies of subinvariants, see Tables 6, 6, 6. We see that frequencies for bivector and sextivector (and pseudoscalar for $A_{8}$) subinvariants for $A_{8}$ and $D_{8}$ are halved, meaning that half of these subinvariants differ from the other half by a minus sign. At the same time, scalar and quadrivector subinvariants are unchanged. The picture is different for $E_{8}$, where bivector, quadrivector and sextivector frequencies change non-trivially under sign identification.
The idea behind these observations is to understand the symmetries of the root system. The frequency of multiplicities for the three algebras should be determined by the permutation of these symmetries. It is rather simple to determine the number of unique invariants and explain the existence of Doublets and Quadruplets for the $A_{8}$ algebra due to its simple Dynkin diagram, but much harder for the $D_{8}$ and $E_{8}$ algebras.
@@ 4. Machine Learning
ML as a field is subdivided into three core categories, : supervised, unsupervised and reinforcement learning. The focus of this study is on the first two. In supervised learning, many-parameter functions are fitted to large datasets of (input, output) pairs. In unsupervised learning there is no respective output data, and traditional data analysis methods for feature extraction and clustering are applied directly to the datasets.
@@ 4.1.1. Binary Classification of Invariants: real vs fake data
The dataset used consists of 3 subsets, one for each of the algebras $A_{8}$/$E_{8}$/$D_{8}$, with 40320 entries and 2304 components in each subset. To further enlarge the amount of data, we generated a ‘fake’ dataset for each of the $A_{8}$/$E_{8}$/$D_{8}$ algebra. In the first iteration, ‘fake’ datasets were generated by constructing empirical distributions for each of the 2304 components from all available invariants for the corresponding algebra and then sampling from this distribution to create 40,000 unique elements in new datasets. One can notice that in the ‘real’ data, the number of zeros in each entry is constant and specific for different algebras 2020201942 for $E_{8}$, 2083 for $D_{8}$ and 1805 for $A_{8}$.. We implemented this feature in fake datasets we used by eliminating fake data entries which do not satisfy this condition.
@@
One should notice however that training NNs to distinguish one of the $A_{8}/D_{8}/E_{8}$ invariants from others and fake invariants using all datasets is invalid. For all three of them, the prediction on test datasets would be 100% accurate because, as described in section 3.2, there are only 128 unique invariants for each of the ADE datasets, leading to a large repetition of them (although with unequal frequencies). Hence, there is a high chance that in a randomly chosen training subset, we would find all 128 unique invariants. This would make the test stage biased, as it likely contains repetitions of the training data. Effectively, these NNs are learning to reproduce this dataset of invariants perfectly, but would not be able to generalise beyond it.
A more meaningful problem is to remove degeneracy in datasets for ‘real’ $A_{8}/D_{8}/E_{8}$ invariants, leaving just 128 elements in each. Then, we can mix original and fake invariants, which ensures that we have some real invariants in the training set and others in the test set only. The proportion of data in training and test datasets was again set to 80% and 20%. However, this makes the whole dataset skewed as there are around 40000 fake invariants and only $3\times 128$ real invariants. We kept data unbalanced in the training set, but in the test set, to make it easier to interpret results, we cut the number of fake invariants to be the same as the number of real invariants, meaning we had $3\times\{128/k$ real and $128/k$ fake} invariants.
In this setup, we tried NN with architectures varying from 1 hidden layer with 32 units to 2 hidden layers with 256 units in each. In all of them, the ReLU activation function was used. During training, we used the Adam optimizer (with a learning rate of 0.001) to optimize the log-loss function. The train and test datasets represented an 80% / 20% split of the total data. The best performance was demonstrated by 1 hidden layer 64, 128 and 256 units NNs with accuracies in the range of 0.90-0.92. From this one might speculate that there should be some relatively simple invariant quantity (similar to the genus of a surface) that was learned by the NNs to distinguish real invariants coming from different algebras and fake invariants.
@@ 4.1.2. Regressing Invariants from Permutations
In this paper, we used dense NNs coded in python where the input is the (one-hot encoded [6]) permutation and the output is the coefficients of the invariants. For clearer results, we first partitioned each dataset for $A_{8},D_{8},E_{8}$ into 9 subdatasets for each of the 9 invariant orders, and then look at the coefficients of both the full invariant and each subinvariant for each order subdataset, i.e., the scalar, bivector, quadrivector, sextivector, and pseudoscalar for each invariant order $\mathop{\mathrm{Inv}}_{0}$-$\mathop{\mathrm{Inv}}_{8}$ (SOCM). The NN model includes four dense layers of 256 units with ReLU activation function. In training, we used the Adam optimizer (with learning rate 0.001) to minimise a mean-squared error loss. 5-fold cross validation was also used, with test data subset being $20\%$ of the full dataset. To calculate accuracy, predictions were rounded to the nearest integer, and a prediction was considered correct if all of its coefficients were predicted correctly after rounding.
Our ML results are summarised in Tables 9 to 9 for the $A_{8},D_{8},E_{8}$ data in the simple root basis. The results show near-perfect prediction of all invariants and subinvariants across all algebras. The trivial scalar invariants are unsurprisingly all learnt perfectly, but in many other cases there is perfect learning also. The lowest performance occurs for the sextivectors, where there is less data to learn from. These results indicate that the NNs are capable of well approximating the complicated algorithm carried out to compute these invariants, as well as accommodating for the basis permutations.
Acc($\mathop{\mathrm{Inv}}_{i}$) Acc($\mathop{\mathrm{Inv}}_{i}^{0}$) Acc($\mathop{\mathrm{Inv}}_{i}^{2}$) Acc($\mathop{\mathrm{Inv}}_{i}^{4}$) Acc($\mathop{\mathrm{Inv}}_{i}^{6}$) Acc($\mathop{\mathrm{Inv}}_{i}^{8}$)
$\mathop{\mathrm{Inv}}_{0}$ $1.0000$ $1.0000$
$\mathop{\mathrm{Inv}}_{1}$ $1.0000$ $1.0000$ $0.9996$
$\mathop{\mathrm{Inv}}_{2}$ $0.9996$ $1.0000$ $0.9999$ $0.9999$
$\mathop{\mathrm{Inv}}_{3}$ $0.9980$ $1.0000$ $0.9973$ $0.9480$ $0.9999$
$\mathop{\mathrm{Inv}}_{4}$ $0.9958$ $1.0000$ $0.9999$ $0.9999$ $0.9117$ $0.9596$
$\mathop{\mathrm{Inv}}_{5}$ $0.9986$ $1.0000$ $0.9995$ $0.9999$ $1.0000$
$\mathop{\mathrm{Inv}}_{6}$ $1.0000$ $1.0000$ $0.9948$ $0.9999$
$\mathop{\mathrm{Inv}}_{7}$ $0.9999$ $1.0000$ $1.0000$
$\mathop{\mathrm{Inv}}_{8}$ $1.0000$ $1.0000$
@@
Acc($\mathop{\mathrm{Inv}}_{i}$) Acc($\mathop{\mathrm{Inv}}_{i}^{0}$) Acc($\mathop{\mathrm{Inv}}_{i}^{2}$) Acc($\mathop{\mathrm{Inv}}_{i}^{4}$) Acc($\mathop{\mathrm{Inv}}_{i}^{6}$) Acc($\mathop{\mathrm{Inv}}_{i}^{8}$)
$\mathop{\mathrm{Inv}}_{0}$ $1.0000$ $1.0000$
$\mathop{\mathrm{Inv}}_{1}$ $1.0000$ $1.0000$ $0.9955$
$\mathop{\mathrm{Inv}}_{2}$ $1.0000$ $1.0000$ $0.9912$ $0.9999$
$\mathop{\mathrm{Inv}}_{3}$ $0.9993$ $1.0000$ $0.9995$ $0.9999$ $1.0000$
$\mathop{\mathrm{Inv}}_{4}$ $0.9995$ $1.0000$ $0.9988$ $0.9891$ $0.9998$ $1.0000$
$\mathop{\mathrm{Inv}}_{5}$ $0.9995$ $1.0000$ $0.9986$ $1.0000$ $1.0000$
$\mathop{\mathrm{Inv}}_{6}$ $1.0000$ $1.0000$ $1.0000$ $1.0000$
$\mathop{\mathrm{Inv}}_{7}$ $1.0000$ $1.0000$ $0.9999$
$\mathop{\mathrm{Inv}}_{8}$ $1.0000$ $1.0000$
@@
Acc($\mathop{\mathrm{Inv}}_{i}$) Acc($\mathop{\mathrm{Inv}}_{i}^{0}$) Acc($\mathop{\mathrm{Inv}}_{i}^{2}$) Acc($\mathop{\mathrm{Inv}}_{i}^{4}$) Acc($\mathop{\mathrm{Inv}}_{i}^{6}$) Acc($\mathop{\mathrm{Inv}}_{i}^{8}$)
$\mathop{\mathrm{Inv}}_{0}$ $1.0000$ $1.0000$
$\mathop{\mathrm{Inv}}_{1}$ $1.0000$ $1.0000$ $1.0000$
$\mathop{\mathrm{Inv}}_{2}$ $0.9999$ $1.0000$ $1.0000$ $1.0000$
$\mathop{\mathrm{Inv}}_{3}$ $0.9994$ $1.0000$ $0.9906$ $0.9999$ $0.9891$
$\mathop{\mathrm{Inv}}_{4}$ $0.9969$ $1.0000$ $1.0000$ $0.9963$ $0.9793$ $0.9223$
$\mathop{\mathrm{Inv}}_{5}$ $0.9994$ $1.0000$ $0.9996$ $0.9999$ $0.9990$
$\mathop{\mathrm{Inv}}_{6}$ $1.0000$ $1.0000$ $1.0000$ $1.0000$
$\mathop{\mathrm{Inv}}_{7}$ $1.0000$ $1.0000$ $0.9998$
$\mathop{\mathrm{Inv}}_{8}$ $1.0000$ $1.0000$